OpenStudy (anonymous):

Suppose that f is even and g is odd. a. show that f o f is even b. show that g of is even.

OpenStudy (freckles):

you are doing the same thing on previous problems are you unsure what the o looking thing between f and f means? it just means f(f(x)) in a an for b you look at g(f(x)) I would do the same thing here Let h(x)=f(f(x)) replace x with -x and try to show h(-x)=h(x) since in a you want even result

OpenStudy (anonymous):

You can show f is even. f o f (-x) = f( f( -x) ) = f ( f (x) ) = f o f ( x)

OpenStudy (anonymous):

g o f (-x) = g(f(-x)) = g( f(x)) = g o f (x) therefore g o f is even